OpenStudy (anonymous):

Given the function f(x) = 3x + 1, evaluate f(a + 1).

OpenStudy (gorv):

put a+1 in place of x f(a+1)=3(a+1)+1 now solve it

OpenStudy (anonymous):

can you solve the problem please

OpenStudy (gorv):

=3a+3+1 =3a+4 that's it